OpenStudy (dominantvampire):

A car is driven east for a distance of 50 km, then north for 30 km, and then in a direction 30 degrees east of north for 25 km. Determine the magnitude and the angle of the car displacement from its starting point

OpenStudy (anonymous):



OpenStudy (dominantvampire):

mmhmm

OpenStudy (anonymous):

vector BC can be resolved into 2 components, one along East and another along North Total horizontal displacement \[x^{\rightarrow}=(50+(25.\sin(30))i\]\[x^{\rightarrow}=62.5i \] total vertical displacement\[y^{\rightarrow}=(30+(25\cos30))j\]\[y^{\rightarrow}=51.65j\] Total displacement \[|R^{\rightarrow}|=\sqrt{|x^{\rightarrow}|^2+|y^{\rightarrow}|^2}\]